OpenStudy (anonymous):

Find dy/du, du/dx, and dy/dx when y and u are defined as follows. y = √u u = 7x - 6x2

OpenStudy (anonymous):

must derive y= sqrt. of u and u=7x-6x2 or is that 6x^2?

OpenStudy (anonymous):

@DarlaD so what did you get when you derive y and u?

OpenStudy (anonymous):

uhmm.. pls. reply? :D can you derive y an u?

OpenStudy (anonymous):

its 6x^2

OpenStudy (anonymous):

okay so what did u get for dy/du and the rest? :)

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

now that we already get dy and du... what is dy/du? :)

OpenStudy (anonymous):

1/2u^(-1/2)/7-12x but how do you reduce that

OpenStudy (anonymous):

sorry my bad wait

OpenStudy (anonymous):

i mean when you derive y it is dy/du so y= sqrt. of u \[\frac{ dy }{ du }= \frac{ 1 }{ 2 }u ^{\frac{ -1 }{ 2 }}\] u = 7x-6x^2 \[\frac{ du }{ dx }= 7 - 12x\] now using the chain rule: dy/dx = dy/du * du/dx \[\frac{ dy }{ dx }= (\frac{ 1 }{ 2 }u ^{\frac{ -1 }{ 2 }}) * (7 - 12x)\]

OpenStudy (anonymous):

since u = 7x - 6x^2 substitute that to \[\frac{ 1 }{ 2 }u ^{\frac{ -1 }{ 2 }}\] it'll be \[\frac{ 1 }{ 2 }(7x-6x^2)^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

Is there a different form Is should put it in? the answer still isn't right

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= (\frac{ 1 }{ 2 }(7x-6x^2)^{\frac{ -1 }{ 2 }})*(7-12x)\]

OpenStudy (anonymous):

what do you mean? :) we're not yet getting th efinal answer :)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So is that not the final answer?

OpenStudy (anonymous):

not yet... wait a second...

OpenStudy (anonymous):